OpenStudy (anonymous):

Let f(x) = - 4x +5, find F(1over2) =

OpenStudy (anonymous):

Change x to 1/2. You will have:\[f(\frac{ 1 }{ 2 }) = -4(\frac{ 1 }{ 2 }) + 5 = \frac{ -4 }{ 2 } + 5 = -2 + 5 = 3\]